• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name xqian

    Log entry time 23:53:33 on November 9,2008

    Entry number 247876

    keyword=Swing Day Shift Summary

    16:00 pm There is still no beam, we are waiting for the beam are we are still taking the cosmic data.
    16:50 We asked for a controlled access
    17:20 We are going back to Beam permit.
    18:30 We start have the tune beam
    18:50 We start have CW beam
    Start with BeO target raster off 2uA
    beam position is good

    Then we moved to the polarized 3He target with 4x4 MCC unit raster on. We checked the rates and the raster size using spot++

    We are going to the 8uA production!

    The VDC is slowly recovering
    RUN 3124 Polarized 3He start 19:19 pm 19:40 pm 1M events

    The timing are "screwed up" for the left HRS in the RUN 3124.

    RUN 3125 Polarized 3He start 19:50pm -> 1M
    RUN 3126 20:23 -> 1M
    RUN 3128 8:55pm -> 1M
    RUN 3129 21:28pm some problems with the BigBite ROCs, deadtime is huge.

    At 23:00 brad took a controlled access to power cycle the ROCs. The problem is solved.

    23:35 moved target to BeO then back to polarized 3He

    RUN3143 11:45pm Polarized 3HE